Output. TotalCost Price. AveageVariableCost. AverageTotalCost. MargianlCost. Profit.

0 $1,000 $500 ? ? ? ?

1 $1,200 $500 ? ? ? ?

2 $1,350 $500 ? ? ? ?

3 $1,550 $500 ? ? ? ?

4 $1,900 $500 ? ? ? ?

5 $2,300 $500 ? ? ? ?

6 $2,750 $500 ? ? ? ?

7 $3,250 $500 ? ? ? ?

8 $3,800 $500 ? ? ? ?

9 $4,400 $500 ? ? ? ?

10 $5,150 $500 ? ? ? ?




a. Complete the above table. What is the firm’s fixed cost? How can you tell? Hint: if the firm is not producing output, it still must pay its fixed cost but incurs no variable cost.




1
Expert's answer
2022-02-18T08:43:58-0500

i) Average Total Cost = Total cost price/output

For output 1,           = 1200/1 = 1200

The rest is calculated filled as shown in the table below:



ii) Marginal cost is calculated as: by dividing change in costs by change in quantity, and the result of fixed costs for items already produced and variable costs that still need to be accounted for.

Marginal cost = current total cost – preceding total cost

Marginal cost for quantity 1 is 1200 – 1000 = 200

The rest is calculated and filled in the table below:



What is the firm’s fixed cost? How can you tell?

The company's fixed cost is $1000. At zero production there exist a cost of $ 1000 which is the fixed cost. The variable cost exists only where there are quantities of goods produced.